OpenStudy (anonymous):

someone please help The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 135 grams of a radioactive isotope, how much will be left after 4 half-lives?

OpenStudy (welshfella):

after first half life it will have a mass of 1/2 * 135 - etc for another 3 half lives

OpenStudy (anonymous):

oay am i just dividing

OpenStudy (anonymous):

the ANSWER WAS 8 ... NEW QUESTION An amount of $24,000 is borrowed for 14 years at 6.5% interest, compounded annually. If the loan is paid in full at the end of that period, how much must be paid back?

OpenStudy (anonymous):

@songgirl1234 @butterflydreamer @iGreen

OpenStudy (anonymous):

@horsegirl325 @Ineedhelpfast12 @barce_messi10

OpenStudy (anonymous):

135/2 =67.5 1 half 67.5/2=33.75 2 33.75/2=16.875 3 16.875/2 =8.4 375 grams

OpenStudy (anonymous):

I already solved that one , the answer was just 8 , it;s the next question but i thank u so much

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

can u help me with the last one

OpenStudy (anonymous):

Compound Interest Formula A= P(1+r/n)^(nt)

OpenStudy (anonymous):

okay